Convert 3.25x105 grams Pb into molecules


1
Expert's answer
2021-02-22T05:56:48-0500

Solution:

Moles of Pb = Mass of Pb / Molar mass of Pb

The molar mass of Pb is 207.2 g mol-1.

Hence,

Moles of Pb = 3.25×105 g / 207.2 g mol-1 = 1568.533 mol

Moles of Pb = 1568.533 mol


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Pb molecules = (1568.533 mol of Pb) × (6.022×1023 molecules / 1 mol) = 9.446×1026 Pb molecules

9.446×1026 Pb molecules = 9.45×1026 Pb molecules


Answer: 9.45×1026 Pb molecules are present in 3.25x10grams Pb.